# Uniform Norm Error Estimates for 2D Turning Point Problem

### v1

@@ 4 Estimates for convection and reaction terms involving discrete Green’s function
$\|u-Iu\|_{L^{\infty}(\Omega_{f})}\leq CN^{-2}\log^{2}N,\quad\quad\|u-Iu\|_{L^{\infty}(\Omega_{c})}\leq CN^{-2}.$
Lemma 5.1 When $\beta\geq 2/3$ and the source point $(x_{m},y_{n})$ lies in $\Omega_{c}\cup\Omega_{f,x}$, then $(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,x}}\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$
Proof. We begin by decomposing the solution $u$ into its regular and singular parts as $u=R+F_{1}+F_{2}+F_{12}$. Consider the source point $(x_{m},y_{n})\in\Omega_{c}$. Hence,
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,x}}=$ $(IR-R,b_{1}G_{mn,x})_{\Omega_{f,x}}+(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,x}}$
$+(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,x}}+(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,x}}$ (4.1)
Consider the first term of the R.H.S. of (4.1) $(IR-R,b_{1}G_{mn,x})_{\Omega_{f,x}}$.
$(IR-R,b_{1}G_{mn,x})_{\Omega_{f,x}}$ $\leq\|IR-R\|_{\Omega_{f,x}}\|b_{1}G_{mn,x}\|_{\Omega_{f,x}}$
@@
$\leq\frac{C}{N^{2}}\sqrt{\varepsilon}\log^{1/2}(1/\varepsilon)\;\|G_{mn,x}\|_{\Omega_{f,x}}$
$\leq\frac{C}{N^{2}}\log^{1/2}(1/\varepsilon)\;\|G_{mn}\|_{1,\varepsilon}\leq\frac{C}{N}\log^{1/2}(1/\varepsilon).$ (4.2)
Consider the second term of the R.H.S. of (4.1) $(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,x}}$.
$(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,x}}$ $\leq\|IF_{1}-F_{1}\|_{\Omega_{f,x}}\|b_{1}G_{mn,x}\|_{\Omega_{f,x}}$
@@
$\leq C\|IF_{1}-F_{1}\|_{L^{\infty}(\Omega_{f,x})}\sqrt{\varepsilon}\log^{1/2}(1/\varepsilon)\;\|G_{mn,x}\|_{\Omega_{f,x}}$
@@
$\leq C\|IF_{1}-F_{1}\|_{L^{\infty}(\Omega_{f,x})}\log^{1/2}(1/\varepsilon)\;\|N$
$\leq CN[h_{x}^{2}\|F_{1,xx}\|_{L^{\infty}(\Omega_{f,x})}+h_{x}h_{y}\|F_{1,xy}\|_{L^{\infty}(\Omega_{f,x})}+h_{y}^{2}\|F_{1,yy}\|_{L^{\infty}(\Omega_{f,x})}]$
$\leq CN\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N^{2}}\frac{1}{\varepsilon^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon)}{N^{2}}\frac{1}{\varepsilon}+\frac{1}{N^{2}}\right]\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$ (4.3)
Next consider the third term of the R.H.S. of (4.1) $(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,x}}$.
$(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,x}}$ $\leq\|IF_{2}-F_{2}\|_{\Omega_{f,x}}\|b_{1}G_{mn,x}\|_{\Omega_{f,x}}$
@@
$\leq C\|IF_{2}-F_{2}\|_{L^{\infty}(\Omega_{f,x})}\sqrt{\varepsilon}\log^{1/2}(1/\varepsilon)\;\|G_{mn,x}\|_{\Omega_{f,x}}$
@@
$\leq CN\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon)}{N^{2}}\frac{1}{\varepsilon^{1/2}}\right]$
$\quad\quad\quad\quad\quad+CN\left[\frac{1}{N^{2}\varepsilon}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}\right].$ (4.4)
Consider
$\frac{1}{\varepsilon}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}\leq\frac{1}{\varepsilon}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}\|_{L^{\infty}(\Omega_{f,x})}+\frac{1}{\varepsilon}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}$ (4.5)
Note that $\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}\|_{L^{\infty}(\Omega_{f,x})}=\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}\|_{L^{\infty}(-1+\lambda_{y},1-\lambda_{y})}$.
Also let $\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}\|_{L^{\infty}(-1+\lambda_{y},1-\lambda_{y})}=e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y_{0})}$ where $y_{0}\in[-1+\lambda_{y},1-\lambda_{y}]$.
When $y_{0}\in[-1+\lambda_{y},1-\lambda_{y}]$, then
$y_{0}$ $\geq-1+\lambda_{y}\implies 1+y_{0}\geq\lambda_{y}$
$\implies\frac{-\beta}{\sqrt{\varepsilon}}(1+y_{0})$ $\leq\frac{-\beta}{\sqrt{\varepsilon}}\lambda_{y}\implies e^{\frac{-\beta}{\sqrt{\varepsilon}}(1+y_{0})}\leq e^{\frac{-\beta}{\sqrt{\varepsilon}}\lambda_{y}}\leq\varepsilon^{\frac{3\beta}{2}}.$ (4.6)
Similarly when $y_{0}\in[-1+\lambda_{y},1-\lambda_{y}]$, then
$y_{0}$ $\leq 1-\lambda_{y}\implies 1-y_{0}\geq\lambda_{y}$
$\implies\frac{-\beta}{\sqrt{\varepsilon}}(1-y_{0})$ $\leq\frac{-\beta}{\sqrt{\varepsilon}}\lambda_{y}\implies e^{\frac{-\beta}{\sqrt{\varepsilon}}(1-y_{0})}\leq e^{\frac{-\beta}{\sqrt{\varepsilon}}\lambda_{y}}\leq\varepsilon^{\frac{3\beta}{2}}.$ (4.7)
also for $x$, when $x\in[\lambda_{x},1]$
$e^{\frac{-\alpha x}{\varepsilon}}$ $\leq e^{\frac{-\alpha\lambda_{x}}{\varepsilon}}=\varepsilon^{\alpha}.$ (4.8)
Combining (4.4 - 4.8), we obtain
$(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,x}}$ $\leq CN\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon)}{N^{2}}\frac{1}{\varepsilon^{1/2}}+\frac{\varepsilon^{\frac{3\beta}{2}-1}}{N^{2}}\right]$
$\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$ (4.9)
Next consider the fourth term of the R.H.S. of (4.1) $(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,x}}$.
$(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,x}}$ $\leq\|IF_{12}-F_{12}\|_{\Omega_{f,x}}\|b_{1}G_{mn,x}\|_{\Omega_{f,x}}$
@@
$\leq CN[h_{x}^{2}\|F_{12,xx}\|_{L^{\infty}(\Omega_{f,x})}+h_{x}h_{y}\|F_{12,xy}\|_{L^{\infty}(\Omega_{f,x})}$
$\quad\quad\quad\quad\quad+h_{y}^{2}\|F_{12,yy}\|_{L^{\infty}(\Omega_{f,x})}]$
$\leq\frac{C}{N}\left[\frac{\varepsilon^{2}\log^{2}(\frac{1}{\varepsilon})}{\varepsilon^{2}}+\|e^{\frac{-\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{f,x})}\frac{h_{x}}{\varepsilon^{3/2}}\right]$
$\quad\quad\quad\quad\quad+CN\left[\frac{1}{N^{2}\varepsilon}\|e^{\frac{-\alpha x}{\varepsilon}}(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)})\|_{L^{\infty}(\Omega_{f,x})}\right]$
$\leq C\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N}+\frac{\varepsilon\log(1/\varepsilon)}{N}\frac{\varepsilon^{\frac{3\beta}{2}}}{\varepsilon^{3/2}}+\frac{\varepsilon^{\frac{3\beta}{2}-1}}{N}\right]\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$ (4.10)
Hence when source point $(x_{m},y_{n})\in\Omega_{c}$ then we have
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,x}}\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$
Similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{f,x}$, then we have
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,x}}\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$
@@
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,x}}\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$
Lemma 5.2 When $\alpha>2$ and the source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, then $(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}\leq C\frac{\log^{2}(1/\varepsilon)}{N}.$
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{c}$ and consider $(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}$ but $u=R+F_{1}+F_{2}+F_{12}.$ Hence,
$(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}=$ $(IR-R,b_{1}G_{mn,x})_{\Omega_{f,y}}+(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,y}}$
$+$ $(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,y}}+(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,y}}$ (4.11)
Consider the second term of the R.H.S. of (4.11) $(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,y}}$.
$(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,y}}$ $\leq\|IF_{1}-F_{1}\|_{\Omega_{f,y}}\|b_{1}G_{mn,x}\|_{\Omega_{f,y}}$
@@
$\leq CN^{2}\varepsilon^{1/4}[h_{x}^{2}\|F_{1,xx}\|_{L^{\infty}(\Omega_{f,y})}+h_{x}h_{y}\|F_{1,xy}\|_{L^{\infty}(\Omega_{f,y})}$
$\quad+h_{y}^{2}\|F_{1,yy}\|_{L^{\infty}(\Omega_{f,y})}]$
$\leq CN^{2}\varepsilon^{1/4}\left[\frac{1}{N^{2}\varepsilon^{2}}+\frac{\sqrt{\varepsilon}\log(1/\varepsilon^{3/2})}{\varepsilon N^{2}}+\varepsilon\log^{2}(1/\varepsilon^{3/2})\right]\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,y})}$
$\leq\frac{C}{N}\quad\text{when}\;\alpha>2.$ (4.12)
Next consider the third term of the R.H.S. of (4.11) $(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,y}}$.
$(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,y}}$ $\leq\|IF_{2}-F_{2}\|_{\Omega_{f,y}}\|b_{1}G_{mn,x}\|_{\Omega_{f,y}}$
@@
$\leq CN^{2}\varepsilon^{1/4}[h_{x}^{2}\|F_{2,xx}\|_{L^{\infty}(\Omega_{f,y})}+h_{x}h_{y}\|F_{2,xy}\|_{L^{\infty}(\Omega_{f,y})}$
@@
$\leq CN^{2}\varepsilon^{1/4}\left[\frac{1}{N^{2}}+\frac{\sqrt{}{\varepsilon}\log(1/\varepsilon^{3/2})}{N^{2}}\frac{1}{\varepsilon^{1/2}}+\frac{\varepsilon\log^{2}{(1/\varepsilon^{3/2})}}{\varepsilon N^{2}}\right]$
$\leq\frac{C\log^{2}({1/\varepsilon^{3/2}})}{N}.$ (4.13)
Next consider the fourth term of the R.H.S. of (4.11) $(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,y}}$.
$(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,y}}$ $\leq\|IF_{12}-F_{12}\|_{\Omega_{f,y}}\|b_{1}G_{mn,x}\|_{\Omega_{f,y}}$

### v2

@@ 5 Estimates related to convection and reaction terms
$\|u-Iu\|_{L^{\infty}(\Omega_{f})}\leq CN^{-2}\log^{2}(1/\varepsilon^{3/2}),\quad\quad\|u-Iu\|_{L^{\infty}(\Omega_{c})}\leq CN^{-2}.$
Lemma 5.2
If the source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, then for $\beta\geq 2/3$ following estimate hold
$|(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,x}}|\leq C\frac{(\log(1/\varepsilon))^{2}}{N}.$
Proof. We begin by decomposing the solution $u$ into its regular and singular parts as $u=R+F_{1}+F_{2}+F_{12}$. Consider the source point $(x_{m},y_{n})\in\Omega_{f,x}$. Hence,
$|(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,x}}|\leq$ $|(IR-R,b_{1}G_{mn,x})_{\Omega_{f,x}}|+|(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,x}}|$
$+|(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,x}}|+|(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,x}}|.$ (5.1)
Consider the first term of the R.H.S. of (5.1) is $|(IR-R,b_{1}G_{mn,x})_{\Omega_{f,x}}|$. By using Theorem 5.1 and furthermore, we obtain
$|(IR-R,b_{1}G_{mn,x})_{\Omega_{f,x}}|$ $\leq\|IR-R\|_{\Omega_{f,x}}\|b_{1}G_{mn,x}\|_{\Omega_{f,x}}$
@@
$\leq\frac{C}{N^{2}}\log^{1/2}(1/\varepsilon)\;\|G_{mn}\|_{1,\varepsilon}\leq\frac{C}{N^{2}}\log^{1/2}(1/\varepsilon)\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq\frac{C}{N^{3/2}}\log(1/\varepsilon).$ (5.2)
Consider the second term of the R.H.S. of (5.1) is $|(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,x}}|$. By using Theorem 1.2, Theorem 4.1, Corollary 3.3.1 and furthermore, we obtain
$|(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,x}}|$ $\leq\|IF_{1}-F_{1}\|_{\Omega_{f,x}}\|b_{1}G_{mn,x}\|_{\Omega_{f,x}}$
@@
$\leq C\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}[h_{x}^{2}\|F_{1,xx}\|_{L^{\infty}(\Omega_{f,x})}+h_{x}h_{y}\|F_{1,xy}\|_{L^{\infty}(\Omega_{f,x})}$
$\quad\quad\quad\quad\quad\quad\quad+h_{y}^{2}\|F_{1,yy}\|_{L^{\infty}(\Omega_{f,x})}]$
$\leq C\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{\varepsilon^{2}N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon)}{N^{2}\varepsilon}+\frac{1}{N^{2}}\right]$
$\leq C\frac{(\log(1/\varepsilon))^{5/2}}{N^{3/2}}.$ (5.3)
Next consider the third term of the R.H.S. of (5.1) is $|(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,x}}|$. By using Theorem 1.2, Theorem 4.1, Lemma 3.4 and furthermore, we obtain
$|(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,x}}|$ $\leq\|IF_{2}-F_{2}\|_{\Omega_{f,x}}\|b_{1}G_{mn,x}\|_{\Omega_{f,x}}$
@@
$\leq C\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon)}{N^{2}}\frac{1}{\varepsilon^{1/2}}\right]$
$\quad\quad\quad\quad\quad+C\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}\left[\frac{1}{N^{2}\varepsilon}\|e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\|_{L^{\infty}(\Omega_{f,x})}\right]$
$\leq C\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)}{N^{2}}+\frac{\varepsilon\log^{2}(1/\varepsilon)}{N^{2}}\frac{1}{\varepsilon^{1/2}}+\frac{\varepsilon^{\frac{3\beta}{2}-1}}{N^{2}}\right]$
$\leq C\frac{(\log(1/\varepsilon))^{5/2}}{N^{3/2}}.$ (5.4)
Next consider the fourth term of the R.H.S. of (5.1) is $|(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,x}}|$. By using Theorem 1.2, Theorem 4.1, Corollary 3.3.1, Lemma 3.4 and furthermore, we obtain
$|(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,x}}|$ $\leq\|IF_{12}-F_{12}\|_{\Omega_{f,x}}\|b_{1}G_{mn,x}\|_{\Omega_{f,x}}$
@@
$\leq C[h_{x}^{2}\|F_{12,xx}\|_{L^{\infty}(\Omega_{f,x})}+h_{x}h_{y}\|F_{12,xy}\|_{L^{\infty}(\Omega_{f,x})}$
$\quad\quad\quad\quad\quad+h_{y}^{2}\|F_{12,yy}\|_{L^{\infty}(\Omega_{f,x})}]\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}$
$\leq C\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}\left[h_{x}^{2}+\frac{h_{x}h_{y}}{\varepsilon^{3/2}}+\frac{h_{y}^{2}}{\varepsilon}\right]\varepsilon^{\frac{3\beta}{2}}$
$\leq C\sqrt{N\log\left(\frac{1}{\varepsilon}\right)}\left[\frac{\varepsilon^{2}\log^{2}(1/\varepsilon)\varepsilon^{\frac{3\beta}{2}}}{N^{2}}+\frac{\varepsilon\log(1/\varepsilon)}{N^{2}}\frac{\varepsilon^{\frac{3\beta}{2}}}{\varepsilon^{3/2}}+\frac{\varepsilon^{\frac{3\beta}{2}-1}}{N^{2}}\right]$
$\leq C\frac{(\log(1/\varepsilon))^{5/2}}{N^{3/2}}.$ (5.5)
Hence when source point $(x_{m},y_{n})\in\Omega_{f,x}$, then we have
$|(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,x}}|\leq C\frac{(\log(1/\varepsilon))^{5/2}}{N^{3/2}}.$
Similar arguments we can show that when source point $(x_{m},y_{n})\in\Omega_{c}$, then we have
$|(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,x}}|\leq C\frac{(\log(1/\varepsilon))^{2}}{N}.$
@@
$|(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,x}}|\leq C\frac{(\log(1/\varepsilon))^{2}}{N}.$
Lemma 5.3
If the source point $(x_{m},y_{n})\in\Omega_{c}\cup\Omega_{f,x}$, then for $\alpha\geq 2$ following estimate hold
$|(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}|\leq C\frac{(\log(1/{\varepsilon}))^{1/2}}{N}.$
Proof. First we take the source point $(x_{m},y_{n})\in\Omega_{f,x}$ and consider $|(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}|$ and decompose $u$ as $u=R+F_{1}+F_{2}+F_{12}.$ Hence,
$|(Iu-u,b_{1}G_{mn,x})_{\Omega_{f,y}}|\leq$ $|(IR-R,b_{1}G_{mn,x})_{\Omega_{f,y}}|+|(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,y}}|$
$+$ $|(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,y}}|+|(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,y}}|$ (5.6)
Consider the second term of the R.H.S. of (5.6) is $|(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,y}}|$. By using inverse inequality, Theorem 1.2, Theorem 4.1, Lemma 3.3 and furthermore, we obtain
$|(IF_{1}-F_{1},b_{1}G_{mn,x})_{\Omega_{f,y}}|$ $\leq\|IF_{1}-F_{1}\|_{\Omega_{f,y}}\|b_{1}G_{mn,x}\|_{\Omega_{f,y}}$
@@
$\leq CN\varepsilon^{1/4}[h_{x}^{2}\|F_{1,xx}\|_{L^{\infty}(\Omega_{f,y})}+h_{x}h_{y}\|F_{1,xy}\|_{L^{\infty}(\Omega_{f,y})}$
$\quad+h_{y}^{2}\|F_{1,yy}\|_{L^{\infty}(\Omega_{f,y})}](N\log(1/{\varepsilon}))^{1/2}$
$\leq CN(N\log(1/{\varepsilon}))^{1/2}\varepsilon^{1/4}\left[\frac{1}{N^{2}\varepsilon^{2}}+\frac{h_{x}h_{y}}{\varepsilon}+h_{y}^{2}\right]\|e^{\frac{-\alpha x}{\varepsilon}}\|_{L^{\infty}(\Omega_{f,y})}$
$\leq\frac{C\varepsilon^{1/4}(\log(1/{\varepsilon}))^{1/2}}{\varepsilon^{2}N^{1/2}}\varepsilon^{\alpha}\leq\frac{C(\log(1/{\varepsilon}))^{1/2}}{N}\quad\text{when}\penalty 10000\ \alpha\geq 2.$ (5.7)
Next consider the third term of the R.H.S. of (5.6) is $|(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,y}}|$. By using inverse inequality, Theorem 1.2, Theorem 4.1, Corollary 3.4.1 and furthermore, we obtain
$|(IF_{2}-F_{2},b_{1}G_{mn,x})_{\Omega_{f,y}}|$ $\leq\|IF_{2}-F_{2}\|_{\Omega_{f,y}}\|b_{1}G_{mn,x}\|_{\Omega_{f,y}}$
@@
$\leq CN(N\log(1/{\varepsilon}))^{1/2}\varepsilon^{1/4}[h_{x}^{2}\|F_{2,xx}\|_{L^{\infty}(\Omega_{f,y})}+h_{x}h_{y}\|F_{2,xy}\|_{L^{\infty}(\Omega_{f,y})}$
@@
$\leq CN(N\log(1/{\varepsilon}))^{1/2}\varepsilon^{1/4}\left[\frac{1}{N^{2}}+\frac{\sqrt{}{\varepsilon}\log(1/\varepsilon^{3/2})}{\varepsilon^{1/2}N^{2}}+\frac{\varepsilon\log^{2}{(1/\varepsilon^{3/2})}}{\varepsilon N^{2}}\right]$
$\leq\frac{C\varepsilon^{1/4}(\log(1/{\varepsilon}))^{1/2}}{N^{1/2}}\leq\frac{C(\log(1/{\varepsilon}))^{1/2}}{N}.$ (5.8)
Next consider the fourth term of the R.H.S. of (5.6) is $|(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,y}}|$. By using Theorem 1.2, Theorem 4.1, Lemma 3.3, Corollary 3.4.1 and furthermore, we obtain
$|(IF_{12}-F_{12},b_{1}G_{mn,x})_{\Omega_{f,y}}|$ $\leq\|IF_{12}-F_{12}\|_{\Omega_{f,y}}\|b_{1}G_{mn,x}\|_{\Omega_{f,y}}$